PI: Edwards

Session: June 1996 (session 2)
Frequency: 6 cm
Mode: E
Obs time (UT): 168 EU 10:00(16/06)-15:00(16/06)
Stations: OnWbSmUr
Processor: Soc
EU Mk3 Tapes/St.: 1


To give feedback on this experiment, click here

Experiment Feedback

Onsala (John Conway at Sun Jun 16 18:22:50 EDT 1996) Minor failures
No known problems

Westerbork (foley at Thu Jun 20 05:51:39 EDT 1996) Success
No known problems

JIVE ( at Tue Jul 2 01:55:43 EDT 1996) Minor failures

Urumqi (Lee Suqin at Tue Jul 2 02:03:26 EDT 1996) Minor failures
It was cloudy; clock offset =0.34 /-0.5us; Tsys=313K; effective factor =49.5%; cal temp =67K. Sorry, We forgot to set CAltemp=67.0 in the PRCfile. Because the tape-start-time setting not advance to the beginning time of each scan in the snapfile, there were about 23 seconds data lost in each scan, No phase cal for 6 cm receiver.

Simeiz (Alexandre Volvatch at Tue Jul 2 06:10:55 EDT 1996) Success
no problems.

HartRAO (Jonathan Quick at Wed Jul 17 08:09:31 EDT 1996) Success
Tape only arrived 25 June 1996. No other problems.